# UniFoil: A Universal Dataset of Airfoils in Transitional and Turbulent Regimes for Subsonic and Transonic Flows

### v2

@@ F.1 Neural network architecture
F.2 Training
$\text{loss}_{\text{encoder-decoder}}$ $=\frac{1}{n_{\text{sample}}}\sum_{i}\frac{\norm{C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}}}{\norm{C^{(i)}_{p,\text{real}}}},$ (14)
$\text{loss}_{\text{DNN}}$ $=\frac{1}{n_{\text{sample}}}\sum_{i}\frac{\norm{C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}}}{\norm{C^{(i)}_{p,\text{real}}}},$
where $C_{p,\text{real}}$ is the pressure coefficient from the ground–truth data and $C_{p,\text{pred}}$ is the pressure coefficient from the predicted data. A plot of the training error as shown in Equation 14 is shown in Figure 9. It is observed that the error steadily decreased for both networks and flattened out after 30 epochs for the DNN while still showing a slight but very slow decreasing trend for the convolutional network layer.
F.3 Prediction
Once the training was complete, a plot for the predicted values of pressure coefficient $C_{p,\text{pred}}$ against the ground truth values of $C_{p,\text{real}}$ was made for all the dataset simulations and their reconstructed field values. This plot is shown in Figure 10. In an ideal case, the network would predict with no loss of accuracy. In such a scenario, we would have a one–one mapping of the predictions and ground truth data, yielding a straight line, $C_{p,\text{real}}=C_{p,\text{pred}}$. This is overlaid onto the plot as well for an ideal–case comparison. In the figure, it is observed that the prediction values were appreciably close to the ground truth values, however the slope of the data does not match the slope of the one–one mapping. In this case predicted $C_{p}$ values are greater then the ground truth near the high end of the plot while predicted $C_{p}$ values near the low end of the plot are less than the ground truth. This shows that the predicted values are scaled slightly more then the ground truth data with the effect becoming more prevalent the further the values are from zero.
Figure 10: Predicted and ground truth values for $C_{p}$.
Next, to quantify the errors in the learning process, all of the input data was regenerated for the flow conditions they were simulated for. An $L_{2}$ error difference study between the actual data and the predicted data for each airfoil was then carried out and the number of airfoils quantity was plotted against the $L_{2}$ error for every airfoil in the 400,000 simulations and is shown in Figure 11. The relative $L_{2}$ error is defined as follows
$\frac{\norm{C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}}}{\norm{C^{(i)}_{p,\text{real}}}},$ (15)
Finally, we present a visualization of the ground truth $C_{p}$ fields, reconstructed $C_{p}$ fields and the $L_{2}$ error for those plots for some of the airfoils randomly picked from the dataset in Figure 12. Appreciable comparisons can be seen for the field values between the ground truth and reconstructed $C_{p}$ data fields.
Figure 12: $C_{p}$ field contours from the ground truth data (first row) and reconstructed data (second row). Absolute error field is plotted in the third row.
We evaluate the model’s accuracy in reproducing the pressure coefficient ($C_{p}$) fields of airfoils by computing the $L_{2}$ norm of the error between the predicted and ground truth values. Notably, higher $L_{2}$ norm errors are observed for the 3rd and 4th airfoils in Figure 12, each with an error of approximately 0.131.
To further examine the model’s performance, we plot the absolute error field (Figure 12),
$\left|C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}\right|,$ (16)
which allows visualization of localized discrepancies. As shown in Figure 12, regions of high error are found in the vicinity of shock waves, particularly over the 3rd and 4th airfoils. These errors typically arise because the model struggles to accurately represent the steep gradients in pressure associated with shocks, resulting in localized deviations along the shock front.
@@ F.3 Prediction
To improve model performance and ensure consistency across varying geometries, two normalization strategies were applied. First, both the model outputs and the dataset were normalized to the range $[-1,1]$. Second, we recorded the minimum and maximum $C_{p}$ values from each simulation and appended them to the model’s latent space. This enabled the network to learn and predict the appropriate scaling factors directly from the geometry.
@@
However, this approach introduces sensitivity: minor inaccuracies in the predicted extrema can propagate as large-scale errors across the entire field. This highlights a key limitation—without directly preserving the true range of the pressure coefficient values, the reconstructed fields may systematically deviate from the physical solution. Future work should explore alternative strategies to better preserve the physical fidelity of predicted $C_{p}$ distributions, particularly in the presence of shocks and weakly varying background fields.

### v4

@@ 4.1 Neural network architecture
4.2 Training results
We present a visualization of the ground truth $C_{p}$ fields, reconstructed $C_{p}$ fields and the $L_{2}$ error for those plots for some of the airfoils randomly picked from the dataset in Figure˜5. Appreciable comparisons can be seen for the field values between the ground truth and reconstructed $C_{p}$ data fields.
To further examine the model’s performance, we plot the absolute error field (Figure˜5),
[…]
which allows visualization of localized discrepancies. As shown in Figure˜5, regions of high error are found in the vicinity of shock waves, particularly over the 3rd and 4th airfoils. These errors typically arise because the model struggles to accurately represent the steep gradients in pressure associated with shocks, resulting in localized deviations along the shock front.
@@ 4.2 Training results
To improve model performance and ensure consistency across varying geometries, two normalization strategies were applied. First, both the model outputs and the dataset were normalized to the range [1,1][-1,1]. Second, we recorded the minimum and maximum $C_{p}$ values from each simulation and appended them to the model’s latent space. This enabled the network to learn and predict the appropriate scaling factors directly from the geometry.
@@
However, this approach introduces sensitivity: minor inaccuracies in the predicted extrema can propagate as large-scale errors across the entire field. This highlights a key limitation—without directly preserving the true range of the pressure coefficient values, the reconstructed fields may systematically deviate from the physical solution. Future work should explore alternative strategies to better preserve the physical fidelity of predicted $C_{p}$ distributions, particularly in the presence of shocks and weakly varying background fields. A detailed discussion, on the training and prediction steps, is in Sections˜D.3 and D.4.
Figure 5: $C_{p}$ field contours from the ground truth data (first row) and reconstructed data (second row). Absolute error field is plotted in the third row. This is the absolute error between the pressure coefficient values from ADflow and prediction from neural network, as shown in the first formula from top in the third row. In the third row, the numbers in each sub–figure represent the relative $L_{2}$ norm error as defined in the second formula from top.
5 Conclusion
6 Acknowledgement
Appendix A Geometry sampling
$\mathbf{Y}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\mathrm{T}},$ (3)
$\widetilde{\mathbf{U}}\in\mathbb{R}^{n\times r},$ (4)
$\mathbf{y}=\widetilde{\mathbf{U}}\tilde{\mathbf{y}}$, (5)
The choice of the modal-coefficients is key in generation of the airfoil shape database. Thus, instead of representing each airfoil with the ($x,y$) coordinate values of its profile, we use the reduced few modal coefficients that were used to generate that airfoil in the input data to the neural network. For more details on choice of modal coefficients and how the sampling is done exactly for large number of airfoils, we direct the readers to Li et al. [2019, 2020]. In this manner, the $4,800$ NLF airfoils and $30,000$ FT airfoils were sampled. The code to sampling these airfoils and list of modal coefficients used for the generation of the dataset are detailed in the GitHub repository publicly available in the CC-BY-SA license2.
Appendix B Numerical method
B.1 Governing equations
The turbulence source term is defined as follows
[…]
B.2 Grid generation
We use pyHyp Secco et al. [2021] to generate the volume grid, which uses a hyperbolic volume grid marching scheme to extrude structured surface grids into volume grids. Thus, the grid generation was performed such that the domain is an O-type grid with the airfoil body–fitted to the center of the domain with one cell thickness in the $z$-direction as shown in Figure˜3.
Table 5: Fine grid generation parameters
Description Value
Number of points along airfoil surface till TE 293
Number of TE points 11
Number of points in marching direction 85
First layer thickness $10^{-6}$ c
Distance from the airfoil surface to march the hyperbolic grid $100c$
To maintain the same shape for the input data format, we ensured to maintain the same input parameters for the grid generation. The input parameters are summarized in Table˜5. In the table, “TE” stands for trailing edge and “cc” stands for chord, which was set to a value of 1m in all simulations.
Appendix C Output file contents
As introduced in Section˜3, the dataset comprises three categories of simulations:
  1. 1.
    Fully turbulent simulations for FT airfoils,
  2. 2.
    Fully turbulent simulations for NLF airfoils,
  3. 3.
    Transition simulations for NLF airfoils.
For each of the output files, we describe the contents in the following subsections. All simulation output files follow a consistent naming convention and are provided in CGNS format. The structure and content of these files are summarized in Table˜6. In the following subsections, we discuss the output folders and files content into two sub–categories based on the geometry, the first being the FT–airfoil dataset and second being the NLF airfoil dataset.
Governing equation Airfoil Folder Suffix Freestream File
C.1 FT–airfoil dataset
Turb_Cutout_<1--6>.zip.
Each simulation file follows the naming pattern:
Airfoil_Case_Data_turb.tab .
A visualization of the field data present in the fully–turbulent dataset is shown in Figure˜6. In the figure, we show the Mach number, velocity magnitude and pressure coefficient values. It also showcases shocks on both the suction and pressure sides, with flow separation aft the shock foot. The velocity magnitude $|\mathbf{u}|$ is computed using the formula
$|\mathbf{u}|=\sqrt{u^{2}+v^{2}},$ (11)
C.2 NLF–airfoil dataset
NLF_Airfoils_Fully_Turbulent.zip ,
with each simulation file inside this folder that follows the name:
Airfoil_Case_Data_Trans_lam.tab ,
for both the fully-turbulent and transition datasets of the NLF–airfoils. Each CGNS file, once again, contains the Mach number, pressure coefficient, $C_{p}$, and the velocity vector field, as shown in Figure˜6 for the FT-airfoils.
airfoil_<airfoil_num>_G2_A_L0_case_<case_num> .
Each of these subfolders contains the following files:
Finally, the archive Transi_sup_data_cutout_<1--4>.zip mirrors this folder structure and includes eight supplementary files per case. These provide extended data on the transition characteristics, enabling deeper analysis of the laminar-to-turbulent transition region. A full description of these supplementary files and their contents is available in the accompanying README file on the GitHub repository2.
A visualization of the velocity magnitude field and the field of ratio of turbulent eddy viscosity to the molecular viscosity is shown in Figure˜7. The transition onset points are marked on the airfoil surface using the red–colored arrow on the suction side of the airfoil. The transition to turbulence was only observed for the suction side in this case. The location for transition point was obtained from the “transiLoc.dat” file.
In the $e^{N}$ transition model, the reduced TS wave frequency is given by
$f=\frac{\omega}{Re_{\delta}},$ (12)
$N=\int_{x_{0}}^{x}\alpha_{i}(x,f)\,\text{d}x,$ (13)
${N}_{\text{TS}}=\text{max}_{f}(N(f)).$ (14)
Out of the eight files, the key files are “nfactor_ts.dat” and “transiLoc.dat”. The former has the TS wave factor $N_{\text{TS}}$ and the latter has the transition point location, if detected, on the suction and pressure surfaces. Details on accessing and visualizing these files can be found in the GitHub repository2. Details on the transition characteristics and the factors can be found in the paper by Shi et al. [2020].
Figures of the skin-friction coefficient $C_{f}$ and the $N_{\text{TS}}$ factor plotted against $x/c$ from the nfactor_ts.dat file is shown in Figure˜8, for the airfoil shown in Figure˜7. The $N_{\text{TS}}$ factor for TS waves is computed as shown in Equation˜14. When transition occurs, the skin–friction coefficient starts to abruptly increase. The transition region highlighted in Figure˜8 can be seen for the suction–side of the airfoil where the skin friction coefficient suddenly rises to a value of roughly 0.0043 before it starts to dip. For the $N_{\text{TS}}$ factor, values begin to increase around $x/c$ of 0.2, in–line with the observation from the skin–friction coefficient plot.
Appendix D Preliminary results for ML–Additional information
D.1 Data pre–processing
To ensure a smooth handling of input data, maintaining the same input shape to the neural network and to preserve the grid resolution without loss of data that typically occurs when interpolating onto a uniform grid, the data from the O-grid was reshaped into a neat 2D array. The process of reshaping is highlighted in Figure˜9. We start at the TE and move clockwise along the airfoil surface to loop till we reach back to the TE. This forms the first bottom–most layer of the 2D array. We then move to the next layer and repeat till all the layers are completed.
Since the same tool (pyHyp) and grid generation inputs from Table˜5 were used, we maintained the same number of “O-grid rings” and ordering of points to reshape all input data in a similar manner. Figure˜9 shows this concept in the top row and shows the implementation of the reshaping algorithm for the velocity field in the bottom row.
D.2 Neural network architecture
The model was trained by first preparing the encoder-decoder model and training it on the reshaped pressure coefficient data. Once this model is trained, the encoder is used to convert all of the pressure coefficient data into latent space values which are saved separately. These latent space values are then combined with matching airfoil feature data in order to train the deep neural network. When the deep neural network is done training, the model can be used to predict an airfoils pressure coefficient field by first converting the feature data into the latent space and then passing this latent space data into the decoder. This results in a pressure coefficient field in the square grid format which must finally be reshaped into the O-Grid using the geometry data from the airfoil.
D.3 Training
$\text{loss}_{\text{encoder-decoder}}$ $=\frac{1}{n_{\text{sample}}}\sum_{i}\frac{\norm{C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}}_{2}}{\norm{C^{(i)}_{p,\text{real}}}_{2}},$ (15)
$\text{loss}_{\text{DNN}}$ $=\frac{1}{n_{\text{sample}}}\sum_{i}\frac{\norm{C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}}_{2}}{\norm{C^{(i)}_{p,\text{real}}}_{2}},$
where $C_{p,\text{real}}$ is the pressure coefficient from the ground–truth data and $C_{p,\text{pred}}$ is the pressure coefficient from the predicted data. A plot of the training error as shown in Equation˜15 is shown in Figure˜10. It is observed that the error steadily decreased for both networks and flattened out after 30 epochs for the DNN while still showing a slight but very slow decreasing trend for the convolutional network layer.
D.4 Prediction
Once the training was complete, a plot for the predicted values of pressure coefficient $C_{p,\text{pred}}$ against the ground truth values of $C_{p,\text{real}}$ was made for all the dataset simulations and their reconstructed field values.
This plot is shown in Figure˜11 for a randomly picked airfoil. In an ideal case, the network would predict with no loss of accuracy. In such a scenario, we would have a one–one mapping of the predictions and ground truth data, yielding a straight line, $C_{p,\text{real}}=C_{p,\text{pred}}$. This is overlaid onto the plot as well for an ideal–case comparison. In the figure, it is observed that the prediction values were appreciably close to the ground truth values. In this case predicted $C_{p}$ values are greater then the ground truth near the high end of the plot while predicted $C_{p}$ values near the low end of the plot are less than the ground truth.
Figure 11: Predicted and ground truth values for $C_{p}$.
Next, to quantify the errors in the learning process, all of the input data was regenerated for the flow conditions they were simulated for. An $L_{2}$ norm error difference study between the actual data and the predicted data for each airfoil was then carried out and the number of airfoils quantity was plotted against the $L_{2}$ error for every airfoil in the 400,000 simulations and is shown in Figure˜12. The relative $L_{2}$ norm error is defined as follows
$\frac{\norm{C^{(i)}_{p,\text{real}}-C^{(i)}_{p,\text{pred}}}_{2}}{\norm{C^{(i)}_{p,\text{real}}}_{2}},$ (16)